square root(x+8) = square root(2x+7)-5
this is what i've gotten so far.
1. (x+8)=2x+32-10squareroot(2x+7)
2. -x-24=-10squareroot(2x-7)
3. x^2 - 200x -124=0
4. quadratic equation outcomes are 200.6180898 and -.6180898248.
is that right?

Answer by stanbon(57285)

sqrt(x+8) = sqrt(2x+7) - 5
Square both sides to get:
x+8 = 2x+7 - 10sqrt(2x+7) + 25
10sqrt(2x+7) = x + 24
Square both sides to get:
2x+7 = x^2 + 48x + 576
x^2 + 46x + 569 = 0
----
This quadratic has no Real Number solutions.
-------
Cheers,
Stan H.